# Unified Triplet-Level Hallucination Evaluation for Large Vision-Language Models

### v2

@@ 3.3 Evaluation Pipeline
NLI Judge.
The first strategy is implemented with a natural language inference (NLI) [29] model  [^1] https://huggingface.co/sentence-transformers/all-mpnet-base-v2. Specifically, given an extracted triplet, we first calculate its cosine similarity scores with all triplets in the image scene graph $G$ and only retain those ground truth (GT) triplets with similarity scores greater than 0.5 to refine the information that will be used for the NLI model. If no triplets in $G$ meet this criterion, only the top three GT triplets with the highest similarity scores will be kept, which are then taken as ground truth inputs for the NLI model to make predictions. If the NLI score between the extracted triplet and ground truth triplets is lower than 0.6, suggesting the extracted triplet cannot be induced based on GT triplets, and therefore, resulting in a hallucination.
# Images # Questions # Objects # Relations # Questions / Image # Triplets / Scene Graph
300 1226 1723 618 4.09 19.10
Table 1: Statistics of Tri-HE.
GPT-4 Judge.
The other strategy is via prompting the GPT-4 model, which is a common practice for recent LLM evaluation literature [30]. We prompt GPT-4 to justify whether a previously extracted triplet $(v_{1},e,v_{2})\in G_{\theta}$ can be directly obtained or inferred based on the image scene graph $G$. Note that we do not adopt the GPT-4V judge, since [31] have observed that the text-only GPT-4 judge better aligns with the human preference than GPT-4V. Moreover, if $(v_{1},e,v_{2})$ is recognized as a hallucinated triplet, we further require it to tell whether the hallucination relates to the relation $e$ or objects $v_{1},v_{2}$. Check our prompt for GPT-4 judge in Figure 8  [^2] For both knowledge graph extraction and GPT-4 judge, we use the “gpt-4-1106-preview” model via OpenAI’s API calls with its default inference parameters..
With triplet-level hallucination judgments in hand, we can adopt Eq. (1) and Eq. (2) to calculate the corresponding hallucination rates of $A_{\theta}$, which can be further separated into object hallucination and relation hallucination rates if GPT-4 judge is adopted. Check more detailed results in Table 2.
4 Tri-HE Construction
Following the formulation in §3, in this section, we provide a detailed discussion on how to construct our Tri-HE benchmark for a unified triplet-level evaluation of both hallucinations in LVLMs.
Dataset.
We build Tri-HE upon the GQA dataset [32], which requires LVLMs to conduct commonsense reasoning based on a given image. The provision of scene graph annotations aligns with our formulation of hallucination evaluation at the triplet-level.
Data collection.
However, several scene graphs in GQA are only equipped with parts of the relations among objects appearing in the image, making necessary information required for question answering absent. To address that, we first employ a filtering process to identify images with adequate distinct relations in the scene graph annotation. Specifically, we start by selecting 300 images with more than 5 relations from the corresponding GQA scene graph annotations. To enlarge the question-answer pairs for each GQA image, we utilize GPT-4V  [^3] We use the “gpt-4-vision-preview” model here, as in §5.2. to generate new questions requiring commonsense reasoning based on the input image. By leveraging the prompt provided in Figure 7, we prompt GPT-4V to generate ten questions for each image along with their corresponding answers and relation triplets that illustrate the reasoning process. Afterward, we manually verify the correctness of generated questions, answers, and triplets, while only questions indeed requiring reasoning are maintained. Moreover, the verified triplets are added to the original scene graphs collected in GQA.
Method GPT-4 Judge NLI Judge
@@ 4 Tri-HE Construction
InternLM2 38.83 37.54 18.25 17.50 20.58 20.04 54.41 52.08
Table 2: Comparison on hallucination rates among different LVLMs on Tri-HE. The best results under each column are boldfaced. InternLM2 is short for InternLM-XComposer2 [33]
Statistics.
The overall statistics for Tri-HE are summarized in Table 1. As described in Figure 2, each image in Tri-HE is linked to a scene graph and a set of question-answer pairs that require reasoning, accompanied by ground truth triplet annotations. Note that since the quality of each question in Tri-HE is manually verified, expanding its size requires significant resources and poses challenges. Nonetheless, the number of images and questions in Tri-HE is comparable to existing LVLM hallucination evaluation benchmarks such as [23] and [8]. Furthermore, as demonstrated in §5, Tri-HE is able to produce reliable hallucination evaluation results.
5 Results on the Tri-HE Benchmark
5.1 Evaluated LVLMs
We selected six open-source LVLMs for evaluation, including the LLaVA series [14, 2], MiniGPT-4 [3], InstructBLIP [1], Shikra [15], and InternLM-XComposer2 (abbrev., InternLM2) [33]. For all evaluated LVLMs, we selected the 7B variants to ensure fair comparison. Additionally, we test the recent popular Llama-3.2-Vision-Instruct model (abbrev., LLaMA-3.2) [34] and used its smallest version (11B). The official prompt templates and inference configurations were used for all LVLMs. All experiments are conducted on two Nvidia A100 GPUs.
5.2 Main Result
LVLM comparison.
Table 2 compares hallucination rates of different LVLMs on our Tri-HE benchmark. As can be seen, all the evaluated LVLMs suffer from generating hallucinations with at least 38% hallucination rates. Among these LVLMs, InternLM2 [33] obtains the best overall performances, suggesting that its strategy to train with both text-image and textual-only instruction data simultaneously helps better align its visual encoder and LLM, and thus, reduces its hallucination rates. Moreover, compared to LLaVA [2], Shikra [15] has consistently lower hallucination rates, which is built upon LLaVA’s structure with extra grounding capability introduced, indicating that introducing extra grounding could help LVLMs reduce hallucination. Additionally, LLaMA-3.2 achieves the lowest relation hallucination rates, suggesting that a strong textual backbone can help mitigate relation hallucination. However, it exhibits a weaker ability to accurately identify objects, impacting its object and overall hallucination rates. Since LLaMA-3.2 does not outperform other LVLMs with even more parameters, we do not adopt it in the remaining.
Relation hallucination is more severe.
Except for MiniGPT-4 and LLaMA-3.2, all the LVLMs generate more relation hallucinations than object hallucinations. A possible explanation is that existing LVLMs lack reasoning abilities, which makes them easily confused and mess up the relations among objects. This further suggests that focusing on object hallucination [5] is not enough for a throughout analysis of the LVLM reliability and a unified and comprehensive study like our proposed triplet-level evaluation is necessary.
Evaluation pipeline.
In addition, we observe that GPT-4 judge can provide clearer and more reasonable discrimination between models compared to NLI judge. We provide a more comprehensive investigation into the differences between these two judges later in §5.3. Besides, the evaluation results under both $\text{Hallu}_{\text{I}}$ and $\text{Hallu}_{\text{Q}}$ metrics demonstrate the same trend, proving the robustness of our proposed triplet-level hallucination evaluation setting under different evaluation granularities.
Method LLaVA LLaVA-1.5 MiniGPT-4 InstructBLIP Shikra InternLM2 GPT-4V
@@ 5.2 Main Result
GPT-4 Judge (Sentence) 0.6631 0.5409 0.3669 0.5532 0.5821 0.5998 0.5548
GPT-4 Judge (Triplet) 0.8115 0.6320 0.4283 0.6235 0.6939 0.7169 0.7292
Table 3: Pearson correlation scores among automatic hallucination judgments and human judgments. The best results under each column are boldfaced.
LLaVA LLaVA-1.5 MiniGPT-4 InstructBLIP Shikra InternLM2
@@
Table 4: Relation hallucination rates for the top 20% frequent object pairs of different LVLMs under the GPT-4 judge. Original refers to the results in Table 2.
Performance of GPT-4V.
Except for the open-sourced LVLMs, we also explore the performance of state-of-the-art commercial LVLMs [13, 35]. Due to the experiment budget, here we only evaluate the GPT-4V [13] model with 25 randomly selected samples in Tri-HE. Specifically, we directly prompting GPT-4V to gather its answers and calculate the corresponding hallucination rates, following the same setting in Table 2. We also report the performances of open-sourced LVLMs on the 25 selected images. As can be seen in Figure 1(b), GPT-4V demonstrates its superior capability and outperforms all open-sourced LVLMs. Although its object hallucination rates are slightly higher than InternLM2 since it tends to associate many objects that are not in the image, GPT-4V has significantly lower reasoning hallucination rates than other LVLMs, indicating its strong reasoning ability, and therefore, reports superior overall hallucination rates.
5.3 Analysis
In §3, we propose to measure hallucination on triplet-level and design two automatic hallucination judges. Here, we further illustrate the effectiveness of the triplet-level evaluation setting by studying its correlation with human judgments. To conduct fine-grained hallucination analysis, previous works [36, 37] split a model response into sub-sentences first, on which their hallucination measurements are conducted. We regard this method as a baseline for comparison. Specifically, we sample a subset of 20 images from Tri-HE and invite human annotators to score five-point-scale hallucination rates of the responses of all the LVLMs in §5.1. Check Table 9 for the detailed annotation guidelines. The human annotators achieve a Krippendorff’s alpha score [38] of 0.66, indicating a high inter-agreement.
Results are shown in Table 3. We find that triplet-level hallucination rates have higher correlations with human judgments with both NLI and GPT-4 judges, indicating that identifying hallucination on triplets can lead to a more accurate, human-preferred evaluation for model responses. Moreover, we notice that the GPT-4 judge achieves a higher correlation to human judgments compared to the NLI counterpart, revealing GPT-4’s superior ability to find hallucinations, which is also consistent with our observation in §5.2.
@@ 5.3 Analysis
Investigating relation hallucination with object information.
As concluded from §5.2, existing LVLMs tend to generate both object and relation hallucinations in their replies, while the relation hallucination rates are even higher. Since different LVLMs have pairs of objects $(v_{1},v_{2})$ that they are familiar with (e.g., high-frequency object pairs in the instruction data they are fine-tuned on) and might generate correct relations on these objects easily, we suppose that the relation hallucination problem might mostly be located in less-frequent object pairs. To verify this assumption, we extract all object pairs for each LVLM from their respective $G_{\theta}$ generated from responses on Tri-HE, and rank these pairs based on their frequency  [^4] Details of this process can be seen in Appendix D.. Then, we calculate each LVLM’s relation hallucination rates on their most frequent object pairs.
As in Table 4, all the LVLMs have significantly lower relation hallucination rates on frequent object pairs they are familiar with, suggesting that they know the possible relations among objects and understand how to choose a relation appropriately. Check more analyses regarding the impact of input length on hallucination rates in Appendix A.
Mitigation GPT-4 Judge NLI Judge
@@
General Description 46.50 49.19 54.59 53.03
Triplet Description 44.14 42.96 51.19 47.12
Table 5: Hallucination mitigation results on MiniGPT-4. The best results under each column are boldfaced.
5.4 Hallucination Mitigation
Regarding the reasons of LVLM hallucination, previous studies [36, 39] have identified modality misalignment as a key factor. Inspired by ECSO [40], we propose a training-free LVLM hallucination mitigation method via self-alignment [12, 40, 41]. Specifically, we disable the direct visual access of LVLMs to alleviate hallucinations resulting from modality misalignment and fully leverage instruction-following capabilities of LVLMs’ (text-only) LLM backbone for hallucinations mitigation [24].
Method.
We start by prompting LVLMs to describe the image on awareness of the given question (i.e., General Description). Subsequently, we prevent LVLMs from having direct access to the image to downgrade the text-image LVLMs to text-only LLMs, which are then encouraged to respond only based on the previously generated image description but without access to the image.
As discussed in §5.3, better hallucination evaluation is observed at the triplet level. Thus, we further prompt LVLMs to focus more on identifying objects and inter-object relations (i.e., Triplet Description) as triplets when describing the image. We use the subset in Figure 1(b) with MiniGPT-4, with the prompts and an example in Figure 3.
Results.
As shown in Table 5, our results demonstrate the effectiveness of hallucination mitigation through our proposed prompting strategies. Both general and triplet descriptions achieve a significant hallucination rate reduction while prompting LVLMs to focus on objects and inter-object relations achieves the lowest hallucination rates. Additionally, we conduct ablation studies and compare our method with existing baselines in Appendix B.
6 Conclusion
Limitations
Due to the experiment budget, in Figure 1(b), we evaluate GPT-4V on a randomly selected subset of 25 images; in Table 3, we conduct human annotation on another randomly selected subset of 20 images. Although we do not interfere with the data selection process, there may still exist a few biases in the corresponding results. Also, we do not have the budget to evaluate more LVLMs on Tri-HE. We will try to provide more comprehensive results in future works once we get more experimental budgets.
Appendix
Appendix A More Analyses
Investigating hallucination rates with response length.
Previous studies on LVLM hallucination evaluation suggest that the length of model responses may influence the extent of hallucination [5, 39], as some LVLMs tend to produce shorter, safer outputs. However, directly instructing an LVLM to generate a response of a specific length is challenging. To address this, we instead truncate the responses to the first $K$ tokens and compute hallucination rates, varying $K$ to assess its impact on the results.
Appendix B More on Hallucination Mitigation
Ablation study.
We conducted an ablation study on a general description (i.e., “describe the image”). As shown in Table 6, the combination of both triplet description prompting and disability of visual input contributed to the effectiveness of our hallucination mitigation approach.
Mitigation GPT-4 Judge NLI Judge
w/o Mitigation 45.86 47.44 55.93 54.94
Ours (eyes-close) 46.50 49.19 54.59 53.03
Ours (triplet) 45.65 45.16 59.35 55.57
Ours (triplet+ eyes-close) 44.14 42.96 51.19 47.12
Table 6: Ablation study on MiniGPT-4 [3]. The best results under each column are boldfaced.
Mitigation on LLaVA-1.5.
To provide a comprehensive evaluation, we conducted additional experiments using another LVLM (i.e., LLAVA-1.5) to mitigate hallucinations. The results presented in Table 7 indicate that both the general description and triplet description methods show an improved reduction in hallucinations, further highlighting the effectiveness of our mitigation approach.
Mitigation GPT-4 Judge NLI Judge
w/o Mitigation 30.72 30.17 53.84 52.06
General Description 28.70 29.80 51.40 49.80
Triplet Description 28.39 32.68 48.97 48.40
Table 7: Further experiment conduct on LLaVA-1.5 [14]. The best results under each column are boldfaced.
Baseline Comparison.
Current methods for mitigating hallucinations typically involve retraining, integrating external detection models, and devising decoding strategies. Compared to existing works, our approach is a plug-and-play method that neither requires costly retraining nor relies on external models. To make a more fair comparison, we conducted experiments with LogicCheckGPT [42], a training-free approach that addresses hallucinations through prompting interactions with the help of GPT-3.5. Under the cost consideration, an evaluation was conducted only with the NLI judge. The results indicate that our method outperforms LogicCheckGPT, highlighting its effectiveness in mitigating hallucinations as shown in Table 8.
@@ Appendix B More on Hallucination Mitigation
Table 8: Hallucination mitigation results on LLaVA-1.5 [14] and MiniGPT-4 [3] with baseline comparison. The best results under each column are boldfaced.
Qualitative comparison
with LogicCheckGPT [42] is provided in Figure 5.
Appendix C NLI Threshold Selection
We randomly selected question instances from 10 images and reviewed the set of filtered triplets that were returned. The similarity score threshold was adjusted to 0.5 for the most reasonable returned triplets. These triplets later concatenate together as the ground truth required for generating NLI judgments. In determining if a generated triplet was hallucinated, we further review the NLI judgment results in different thresholds, ultimately deciding on a threshold of 0.6.
Appendix D Object Pairs Extraction and Ranking
In this section, we detailedly describe how we obtain object pairs and their rankings from LVLM responses. Suppose we have all an LVLM’s responses to all questions in Tri-HE, i.e., $G_{\theta}$, we first extract all the object pairs $(v_{1},v_{2})$ from $G_{\theta}$. Then for each object, we replace it by the name of its synset using WordNet to reduce the total types of objects. Afterward, we could calculate the frequency of each object pair and rank them based on their frequency. This ranking will then be used to calculate the first 20% frequent object pairs in Table 4.
Appendix E More on Prompts
E.1 Prompt for triplets extraction with GPT-4

### v3

@@ 4 Tri-HE Construction
5 Evaluation Results
5.1 Evaluated LVLMs
We selected six open-source LVLMs for evaluation, including the LLaVA series (Liu et al., 2023a), MiniGPT-4 (Zhu et al., 2023), InstructBLIP (Dai et al., 2023), Shikra (Chen et al., 2023d), and InternLM-XComposer2 (abbrev., InternLM-X2) (Cai et al., 2024). For all evaluated LVLMs, we selected the 7B variants to ensure fair comparison. Additionally, we test the recent popular Llama-3.2-Vision-Instruct (abbrev., LLaMA-3.2) (MetaAI, 2024a) and use its smallest version (11B). The prompt templates and inference configurations used for LVLMs are detailed in Appendix §A.4 and §B. All experiments are conducted on two Nvidia A100 GPUs.
5.2 Main Result
LVLM comparison.
Table 2 compares hallucination rates of different LVLMs on our Tri-HE benchmark. As can be seen, all the evaluated LVLMs suffer from generating hallucinations with at least 38% hallucination rates. Among these LVLMs, InternLM-X2 (Cai et al., 2024) obtains the best overall performances, suggesting that its strategy to train with both text-image and textual-only instruction data simultaneously helps better align its visual encoder and LLM, and thus, reduces its hallucination rates. Moreover, compared to LLaVA (Liu et al., 2023b), Shikra (Chen et al., 2023d) has consistently lower hallucination rates, which is built upon LLaVA’s structure with extra grounding capability introduced, indicating that introducing extra grounding could help LVLMs reduce hallucination. Additionally, LLaMA-3.2 achieves the lowest relation hallucination rates, suggesting that a strong textual backbone can help mitigate relation hallucination. However, it exhibits a weaker ability to accurately identify objects, impacting its object and overall hallucination rates. Since LLaMA-3.2 does not outperform other LVLMs with even more parameters, we do not adopt it in the remaining experiments for parameter consistency.
Relation hallucination is more severe.
Except for MiniGPT-4 and LLaMA-3.2, all the LVLMs generate more relation hallucinations than object hallucinations. A possible explanation is that existing LVLMs lack reasoning abilities, which makes them easily confused and mess up the relations among objects. This further suggests that focusing on object hallucination (Li et al., 2023e) is not enough for a throughout analysis of the LVLM reliability, and a unified and comprehensive study like our triplet-level evaluation is necessary.
Evaluation pipeline.
In addition, we observe that LLM Judge can provide clearer and more reasonable discrimination between models compared to the NLI judge. We provide a more comprehensive investigation into the differences between these two judges later in §5.3. Besides, the evaluation results under both $\text{Hallu}_{\text{I}}$ and $\text{Hallu}_{\text{Q}}$ metrics demonstrate similar trends, proving the robustness of our proposed triplet-level hallucination evaluation setting under different evaluation granularities.
Method LLM Judge NLI Judge
@@ 5.2 Main Result
InternLM-X2 38.83 37.54 18.25 17.50 20.58 20.04 54.41 52.08
Table 2: Comparison on hallucination rates among different LVLMs on Tri-HE. The best results under each column are boldfaced. InternLM-X2 is short for InternLM-XComposer2 (Cai et al., 2024)
Method LLaVA LLaVA-1.5 MiniGPT-4 InstructBLIP Shikra InternLM-X2 GPT-4V
@@
LLM Judge (Llama-3.3 Sentence) 0.4705 0.4842 0.3617 0.2520 0.4941 0.4366 0.4969
LLM Judge (Llama-3.3 Triplet) 0.5138 0.5262 0.4150 0.4798 0.5311 0.5323 0.5519
LLM Judge (GPT-4 Sentence) 0.6631 0.5409 0.3669 0.5532 0.5821 0.5998 0.5548
LLM Judge (GPT-4 Triplet) 0.8115 0.6320 0.4283 0.6235 0.6939 0.7169 0.7292
Table 3: Pearson correlation scores among automatic hallucination judgments and human judgments. The best results under each column are boldfaced. The LLM Judges utilized are specified in the brackets.
Evaluate Closed-sourced LVLMs.
In addition to evaluating open-sourced LVLMs, we further investigate the performance of closed-sourced LVLMs. Due to limited experimental resources, we specifically evaluate the GPT-4V (OpenAI, 2023) model on a subset of 25 randomly selected images from Tri-HE. Specifically, we prompt GPT-4V to obtain responses to all questions related to these selected images and compute the associated hallucination rates following the steps described in Table 2. For comparison, we also include results from open-sourced LVLMs evaluated on the same set of 25 images. As illustrated in Figure 1(b), GPT-4V clearly demonstrates superior performance, surpassing all open-sourced LVLMs. Although GPT-4V exhibits slightly higher object hallucination rates compared to InternLM-X2—likely because it tends to associate additional objects not present in the image—it achieves notably lower relation hallucination rates due to its stronger reasoning capabilities, resulting in lower overall hallucination rates.
5.3 Analysis
Here, we further illustrate the effectiveness of the triplet-level evaluation setting by studying its correlation with human judgments. To conduct fine-grained hallucination analysis, previous works (Jing et al., 2023; Min et al., 2023) split a model response into sub-sentences first, on which their hallucination measurements are conducted. We regard this method as a baseline for comparison. Specifically, we sample a subset of 20 images from Tri-HE and invite human annotators to score five-point-scale hallucination rates of the responses of all the LVLMs in §5.1 (check Appendix §C for the detailed annotation guidelines). The human annotators achieve a Krippendorff’s alpha score of 0.66, indicating a high inter-agreement.
Results are shown in Table 3. We find that triplet-level hallucination rates have higher correlations with human judgments with both NLI and LLM Judges, indicating that identifying hallucination on triplets can lead to a more accurate, human-preferred evaluation for model responses. Moreover, we notice that the LLM Judges achieves a higher correlation to human judgments compared to the NLI counterpart, revealing LLMs’ superior abilities to find hallucinations, which is also consistent with our observation in §5.2.
LLaVA LLaVA-1.5 MiniGPT-4 InstructBLIP Shikra InternLM-X2
@@ 5.3 Analysis
Table 4: Relation hallucination rates for the top 20% frequent object pairs of different LVLMs under the LLM Judge. Original refers to the results in Table 2.
Applying Different LLMs in LLM Judge.
While GPT-4 allows the LLM Judge to produce reliable hallucination evaluations, the associated API expenses could become large when evaluating a large number of examples. To mitigate potential cost constraints, we also examine whether alternative open-source LLMs can serve effectively in LLM Judge. Specifically, we replace GPT-4 with LLaMA-3.3-70B-Instruct (abbrev., Llama-3.3) (MetaAI, 2024b) and re-evaluate all examples listed in Table 3. As shown, similar to GPT-4, Llama-3.3 consistently achieves higher correlation scores at the triplet-level than at the sentence-level. Furthermore, its Pearson correlation scores with human evaluations, while significantly outperforming those obtained using NLI Judge, remain comparable to GPT-4’s results for certain LVLMs. These findings suggest that open-source LLMs can serve as viable alternatives to GPT-4 in LLM Judge, providing reliable evaluation results under tight budget constraints, thereby further validating the robustness of our proposed LLM Judge.
Investigating relation hallucination with object information.
As concluded from §5.2, existing LVLMs tend to generate both object and relation hallucinations in their replies, while the relation hallucination rates are even higher. Since different LVLMs have pairs of objects $(v_{1},v_{2})$ that they are familiar with (e.g., high-frequency object pairs in the instruction data they are fine-tuned on) and might generate the correct relations on these objects easily, we suppose that the relation hallucination problem might mostly be located in less-frequent object pairs. To verify this assumption, we extract all object pairs for each LVLM from their respective $G_{\theta}$ generated from responses on Tri-HE, and rank these pairs based on their frequency. Then, we calculate each LVLM’s relation hallucination rates on their most frequent object pairs.
To obtain object pairs and their rankings from LVLM responses, suppose that for the targeted LVLM, we have its responses to all questions in Tri-HE, i.e., $G_{\theta}$, we first extract all the object pairs $(v_{1},v_{2})$ from $G_{\theta}$. Then, for each object, we replace it with the name of its synset using WordNet to reduce the total types of objects. Afterward, we could calculate the frequency of each object pair and rank them based on their frequency. This ranking will then be used to calculate the first 20% frequent object pairs in Table 4.
As in Table 4, all the LVLMs have significantly lower relation hallucination rates on frequent object pairs they are familiar with, suggesting that they know the possible relations among objects and understand how to choose a relation appropriately for frequently occurring objects.
Investigating hallucination rates with response length.
Previous studies on LVLM hallucination evaluation suggest that the length of model responses may influence the extent of hallucination (Li et al., 2023e; Zhou et al., 2023), as some LVLMs tend to produce shorter, safer outputs. However, directly instructing an LVLM to generate a response of a specific length is challenging. To address this, we instead truncate the responses to the first $K$ tokens and compute hallucination rates, varying $K$ to assess its impact on the results.
5.4 Hallucination Mitigation
After demonstrating that LVLMs exhibit significant hallucination problems, we further explore potential approaches to reduce both object and relation hallucinations. Prior works (Jing et al., 2023; Zhou et al., 2023; Li et al., 2023c; Gou et al., 2024) have suggested that modality misalignment might be a primary cause behind LVLM hallucinations. Motivated by this claim, we propose a training-free method to mitigate hallucinations by improving modality alignment within LVLMs.
@@ 5.4 Hallucination Mitigation
Method.
Specifically, we propose a two-step strategy. Given an image and its corresponding question, we first prompt the evaluated LVLM to generate a description of the image guided by the given question (General Description in Figure 4). Next, we prompt the same LVLM (in a new version without the image memory) using this generated description to answer the question. Through this approach, we effectively leverage the strong instruction-following capability intrinsic to the LVLM’s LLM backbone, instead of requiring the LVLM to simultaneously comprehend the image and answer the question, thereby reducing hallucinations caused by modality misalignment. Moreover, as indicated in §5.3, triplet-level evaluation is more effective than sentence-level evaluation in assessing hallucinations. Hence, we further explicitly guide LVLMs to concentrate more on identifying objects and their interrelations via triplets when describing images (Triplet Description in Figure 4). We evaluate MiniGPT-4 and LLaVA-1.5 combined with our proposed mitigation approaches using the subset previously employed in Figure 1(b). The corresponding prompts along with an example illustration are shown in Figure 4.
Mitigation LLM Judge NLI Judge
MiniGPT-4 w/o Mitigation 45.86 47.44 55.93 54.94
General Description 46.50 49.19 54.59 53.03
Triplet Description 44.14 42.96 51.19 47.12
LLaVA-1.5 w/o Mitigation 30.72 30.17 53.84 52.06
General Description 28.70 29.80 51.40 49.80
Triplet Description 28.39 32.68 48.97 48.40
Table 5: Hallucination mitigation results. The best results under each column are boldfaced.
Results.
As demonstrated in Table 5, both LVLMs exhibit reduced hallucination rates after applying our two-stage mitigation method, indicating that improved modality alignment effectively alleviates hallucinations. In addition, explicitly prompting LVLMs to emphasize objects and their relationships consistently yields the lowest hallucination rates across most cases, further reinforcing the findings presented in Table 3.
Ablation Study.
We also perform an ablation study on the best-performing “Triplet Description” variant of our mitigation approach to gain deeper insights into the role of each module within our proposed method. Specifically, we compare the Triplet Description (i.e., Triplet+Eyes-Close) results obtained by MiniGPT-4 with two alternative setups:
  1. 1.
    Eyes-Close: This setting is equivalent to General Description. Image access is disabled (i.e., eyes-close) while prompting LVLM to answer the question. It is designed to assess the impact of employing triplet-level descriptions.
  2. 2.
    Triplet: This setting is similar to Triplet Description but allows image accessibility. It incorporates both the original image and the generated triplet-level description simultaneously as inputs. It is designed to examine the effects of modality alignment.
The experimental results are presented in Table 6. As shown, the combined use of triplet-level description and restriction of visual input access leads to the lowest hallucination rates. These findings further validate the design choices made in our mitigation method.
Mitigation LLM Judge NLI Judge
@@ Results.
Eyes-Close 46.50 49.19 54.59 53.03
Triplet 45.65 45.16 59.35 55.57
Triplet+ Eyes-Close 44.14 42.96 51.19 47.12
Table 6: Ablation study on hallucination mitigation. The best results under each column are boldfaced.
@@
Table 7: Hallucination mitigation results on LLaVA-1.5 Liu et al. (2023a) and MiniGPT-4 Zhu et al. (2023) with baseline comparison. The best results under each column are boldfaced.
Baseline Comparison.
Current methods for mitigating hallucinations typically involve retraining, integrating external detection models, and devising decoding strategies. Compared to existing works, our approach is a plug-and-play method that neither requires costly retraining nor relies on external models. To make a more fair comparison, we conducted experiments with LogicCheckGPT Wu et al. (2024a), a training-free approach that addresses hallucinations through prompting interactions with the help of GPT-3.5. Under the cost consideration, an evaluation was conducted only with the NLI judge. The results indicate that our method outperforms LogicCheckGPT, highlighting its effectiveness in mitigating hallucinations as shown in Table 7. Qualitative comparison with LogicCheckGPT (Wu et al., 2024a) is provided in Figure 8.
6 Conclusion
Acknowledgments
This work has been made possible by a Research Impact Fund project (RIF R6003-21) and a General Research Fund project (GRF 16203224) funded by the Research Grants Council (RGC) of the Hong Kong Government.
Appendix A Prompts
A.1 Prompt for triplets extraction with GPT-4